. Prove the following statements (the symbols carry their usual meaning):



(a) 2F1(−n, 1, 1; x) = (1 − x)



n



(b) 2F1(−n, 1, 1; −x) = (1 + x)



n



(c) x 2F1(1, 1, 2; x) = log(1 − x)



(d) x 2F1(1, 1, 2; −x) = log(1 + x)



(e) x 2F1




1
Expert's answer
2022-01-16T07:47:04-0500


Using hypergeometric series identity:


"_2F_1(a,b;b;z)=(1-z)^{-a}"


(a)

"_2F_1(-n,1;1;x)\implies\>z=x,and\>a=-n"


"\therefore_2F_1(-n,1;1;x)=(1-x)^{-(-n)}"

"=(1-x)^n"




(b)

In "_2F_1(-n,1;1;-x),z=-x\>and \>a=-n"



"\implies\>_2F_1(-n,1;1;-x)=[1-(-x)]^{-(-n)}"


"=(1+x)^n"


(c)

Using hypergeometric series identity;


"_2F_1(1,1;2;-z)=\frac{In(1+z)}{z}"


Then in the statement

"(x)_2F_1(1,1;2;x)"


"x=-z\implies\>z=-x"


"\therefore\>_2F_1(1,1;2;x)=\frac{In(1+-x)}{x}"


"\therefore\>(x)_2F_1(1,1;2;x)=In(1-x)"



(d)


In the statement


"(x)_2F_1(1,1;2;-x)"

"x=z"


"\therefore\>_2F_1(1,1;2;-x)=\frac{In(1+x)}{x}"


"\therefore\>(x)_2F_1(1,1;2;-x)=In(1+x)"




(e)

Assuming the question statement is;

"x_2F_1(\frac{1}{2},\frac{1}{2},\frac{3}{2};x^2)=" arc Sin "x"


Using hypergeometric series identity

"2F_1(\frac{1}{2},\frac{1}{2};\frac{3}{2};z^2)=\frac{arc\>Sin\>z}{z}"


Then substituting "z=x" in the identity

"2F_1(\frac{1}{2},\frac{1}{2};\frac{3}{2};x^2)" "=\frac{arc\>Sin\>x}{x}"


"\implies(x)_2F_1(\frac{1}{2},\frac{1}{2},\frac{3}{2},x^2)=arc\>Sin\>x"